OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 5, –1, –7, … ?

OpenStudy (anonymous):

the general term is 5-(n-1)6 so in n=20 5-(19*6)=-109

OpenStudy (anonymous):

you see that your common term for this sequence is -6 so you can write out your arithmetic sequence because you also have the first term which is 5: \[a_{n} = 5 + (n - 1)-6 = 5 -6n + 6 = -6n + 11\]

OpenStudy (anonymous):

Common difference is -6

OpenStudy (anonymous):

So to find the 20th term plug into your formula n for 20.

OpenStudy (anonymous):

Thanks mandolino I missed that your common difference not common term is -6

OpenStudy (anonymous):

nayeaddo, I thunk that you dropped parens in your genral term form

OpenStudy (anonymous):

So you have: \[a_{20} = -6(20) + 11 = -109\]

OpenStudy (anonymous):

Nayeaddo, swear you are a lifesaver thanks again for the third time!!:)

OpenStudy (anonymous):

yeah I wanted to simplify it so I removed the parentheses by multiplying by -6.

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

But would any of you understand the child question because i can't....